the length of a rectangle is 2 feet more than the width. the perimeter of the rectangle is 20 feet. find the length
Answer by actuary(112)

Let l = the length of the rectangle.
Let w = the width of the rectangle.
l = 2 + w.
The perimeter of rectangle is 2*length + 2*width and in this problem the perimeter is 20.
Putting this information into an algebraic equation, you have 2*l+2*w = 20. Now you know that l = 2 + w so you can substitue this expression for "l" into the equation for the perimeter. This gives you, 2*(2+w)+2*w=20.
Use the Distributive Property to simplify the equation so you now have
4+2*w+2*w=20. Combining like terms, you have 4+4*w=20. Since 4 is a factor of each term in the equation, you can divide each term of the equation by 4 and you now have 1 + w = 5. Subtract 1 from each side of the equation and you have w=4. Since l = 2+w, l=6.
Therefore the width of the rectangle is 4 feet and the length is 6 feet.
Have fun studying math.
Larry